The value of a delivery van depreciates at the rate of 15% per year. Show that

annual depreciation follows a geometric progression, and use this progression

to determine the value of the van after three years, if it is purchased for $4500?


1
Expert's answer
2022-03-03T05:42:16-0500

After "n-"th years the value is equal to "4500(1-\frac{15}{100})^n=4500\cdot 0.85^n," and hence annual depreciation follows a geometric progression.

After three years the value is equal to

"4500\cdot 0.85^3\approx 2763.53"